# Pragmatic Inference for Moral Reasoning Acquisition: Generalization via Distributional Semantics

### v2

@@ 5.3 Language Modeling
Although fine-tuning can introduce forgetting in LLMs kirkpatrick2017overcoming; liu2024towards and degrade their general capabilities, liu2025diagnosing emphasize that fine-tuning for moral reasoning, when not accounting for the pragmatic nature of morality, tends to yield higher perplexity. This is also verified in other morality-relevant tasks such as social stereotype mitigation liu2025diagnosingperformancetradeoffmoral. To further demonstrate the benefits of our proposed pragmatic inference methods, we calculate the Perplexity of the optimal fine-tuned checkpoints of baseline approaches and our proposed approach. We use the WikiText dataset merity2016pointer to compute perplexity. Figure 3 reports perplexity results across different training dataset sizes and tasks. Models fine-tuned with our pragmatic inference methods consistently exhibit substantially lower perplexity compared to baseline approaches.
@@ 6 Discussion
Pragmatic Inference for Moral Reasoning. Different from existing approaches that train models to map moral situations with or without moral foundations to moral judgment  hendrycks2020aligning; forbes2020social; jiang2021can; ziems2022moral; jiang2025investigating, this paper proposes several possible bridges that effectively link the two. The bridge is built by textualizing the often unsaid variables (e.g., social norms, metapragmatic evaluations) in pragmatic inference, hence closing the gap between distributional semantics leveraged by LLMs and moral reasoning. Including the inferential steps is found to significantly increase the model performance. The findings are in line with previous studies that develop other pragmatic inferential methods for toxic language detection tasks  sap2020social; zhou-etal-2023-cobra; chen2025pragmatic. In fact, the concept of entextualizing pragmatic inferential process may be widely applicable to other non-logical reasoning tasks, such as the understanding of metaphors, humor, and sarcasm, which LLMs have thus far been found deficient in barattieri_di_sanpietro_pragmatic_2023; ruis_goldilocks_2023. Textualzing the inference processes for various pragmatic tasks also help to reveal the types of pragmatic knowledge that can be taught to LLMs and how they can be incorporated into model training.
Moral Foundations Theory for Moral Reasoning. We should note that it is not always uncontroversial to involve MFs in the pragmatic inference, because: (1) there is criticism that moral foundations lack neurological evidence for its role in moral reasoning and may oversimplify how humans perform moral reasoning haidt2011moral; braicovich2022under; and (2) they cannot capture cultural differences in moralities graham2009liberals; graham2011mapping. While the latter still awaits for examinations with mutilingual datasets, our study, nevertheless, finds the effectiveness of MFs in machine learning reinig2024survey, despite their drawbacks in human reasoning. Additionally, there is also neurological evidence linking human neural responses to moral foundations in the process of moral judgment hopp2023moral. This leaves a further question to future studies as to whether there is a MFs-free inferential pathway that allows LLMs to sufficiently emulate or even precede humans in moral reasoning.

### v3

@@ 5.3 Language Modeling
Although fine-tuning can introduce forgetting in LLMs Kirkpatrick et al. (2017); Liu et al. (2024) and degrade their general capabilities, Liu et al. (2025b) emphasize that fine-tuning for moral reasoning, when not accounting for the pragmatic nature of morality, tends to yield higher perplexity. This is also verified in other morality-relevant tasks such as social stereotype mitigation Liu et al. (2025a). To further demonstrate the benefits of our proposed pragmatic inference methods, we calculate the Perplexity of the optimal fine-tuned checkpoints of baseline approaches and our proposed approach. We use the WikiText dataset Merity et al. (2016) to compute perplexity. Figure 3 reports perplexity results across different training dataset sizes and tasks. Models fine-tuned with our pragmatic inference methods consistently exhibit substantially lower perplexity compared to baseline approaches.
@@ 6 Discussion
Pragmatic Inference for Moral Reasoning. Different from existing approaches that train models to map moral situations with or without moral foundations to moral judgment  Hendrycks et al. (2020); Forbes et al. (2020); Jiang et al. (2021); Ziems et al. (2022); Jiang et al. (2025), this paper proposes several possible bridges that effectively link the two. The bridge is built by textualizing the often unsaid variables (e.g., social norms, metapragmatic evaluations) in pragmatic inference, hence closing the gap between distributional semantics leveraged by LLMs and moral reasoning. Including the inferential steps is found to significantly increase the model performance. The findings are in line with previous studies that develop other pragmatic inferential methods for toxic language detection tasks  Sap et al. (2020); Zhou et al. (2023); Chen and Wang (2025). In fact, the concept of entextualizing pragmatic inferential process may be widely applicable to other non-logical reasoning tasks, such as the understanding of metaphors, humor, and sarcasm, which LLMs have thus far been found deficient in Barattieri di SanPietro et al. (2023); Ruis et al. (2023). Textualzing the inference processes for various pragmatic tasks also help to reveal the types of pragmatic knowledge that can be taught to LLMs and how they can be incorporated into model training.
Moral Foundations Theory for Moral Reasoning. We should note that it is not always uncontroversial to involve MFs in the pragmatic inference, because: (1) there is criticism that moral foundations lack neurological evidence for its role in moral reasoning and may oversimplify how humans perform moral reasoning Haidt and Joseph (2011); Braicovich (2022); and (2) they cannot capture cultural differences in moralities Graham et al. (2009, 2011a). While the latter still awaits for examinations with mutilingual datasets, our study, nevertheless, finds the effectiveness of MFs in machine learning Reinig et al. (2024), despite their drawbacks in human reasoning. Additionally, there is also neurological evidence linking human neural responses to moral foundations in the process of moral judgment Hopp et al. (2023). This leaves a further question to future studies as to whether there is a MFs-free inferential pathway that allows LLMs to sufficiently emulate or even precede humans in moral reasoning.